Use the rule y = -x + 7 to fill in the blank.
If (1, y) is on the graph, then y =

Answer by rapaljer(4671)

This is a LOT easier than it looks! All you have to do is realize that (1,y) means that x=1, and you have to find the value of y!

If y=-x +7, just plug in the value of x=1, and find y:
y=-1+7
y=6

That's it! FINAL ANSWER!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus